OpenStudy (anonymous):

The hypotenuse of a right triangle measures 12 ft. One leg of the triangle measures 7 ft. Find the other leg length. Review the work below. What is incorrect in this student’s work and why can the problem not be solved in this way? • Pythagorean Theorem: a2 + b2 = c2 • 122 + 72 = c2 • 144 + 49 = c2 • 193 = c2 • The other leg is about 13.9 ft.

OpenStudy (anonymous):

because c is the hypotenuse, not one of the legs should be \(7^2+b^2=12^2\)

Parth (parthkohli):

\(\Large 49 + b^{2} = 144 \) \(\Large b^{2} = 95 \)

OpenStudy (anonymous):

by using pythagoras theorm hypotenius^2 = perpendicular^2 + base^2 12^2= 7^2 + perpendicular^2 144=49+ perpendicular^2 \[\sqrt{95}\] 9.746794344